OpenStudy (anonymous):

Find the LCM: x^2-9 and 3x-9

OpenStudy (anonymous):

1) 3(x - 3) 2) 3(x + 3) 3) 3(x -3)(x + 3) 4) (x - 3)(x + 3) are the options

OpenStudy (anonymous):

u must factor both expresions

OpenStudy (raden):

x^2 - 9 = (x + 3)(x - 3) 3x - 9 = 3(x - 3) now, can u determine the common of both ?

OpenStudy (anonymous):

that means what

OpenStudy (anonymous):

x-3

OpenStudy (raden):

nevermind that's to get GCF not LCM :) it is 3(x+3)(X-3)

OpenStudy (anonymous):

remember taht The least common multiple is the smallest whole number that is divisible by each of the numbers.

OpenStudy (anonymous):

ok